Circular No. 7708

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V4740 SAGITTARII = NOVA SAGITTARII 2001 No. 3 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V4740 Sgr has been given to the nova reported on IAUC 7706.

J. D. West, Mulvane, KS, reports the following CCD photometry for V4740 Sgr: Sept. 7.085 UT, V = 6.45 ± 0.05; 7.090, I_{c} = 6.11 ± 0.06; 7.131, V = 6.69 ± 0.05. W. Liller, Viña del Mar, Chile, reports broadband V = 6.87 from a CCD image taken on Sept. 7.0039. West provides the following position measured from one of his images: α = 18^{h}11^{m}45^{s}.98, δ = −30^{o}30′49′′.5 (equinox 2000.0). H. Yamaoka, Kyushu University, reports that the closest object visible on the Digital Sky Survey red plate (limiting mag ∼ 20) is a star of red mag ∼ 17 at approximate position end figures 45^{s}.87, 53′′.3. Visual magnitude estimates: Sept. 6.839 UT, 6.6 (A. Pereira, Cabo da Roca, Portugal); 6.915, 6.6 (Pereira); 7.385, 6.5 (J. Bedient, Honolulu, HI); 7.780, 6.9 (K. Hornoch, Lelekovice, Czech Republic).

XTE J1650–500 On IAUC 7707, third item, line 7, the Mt. John Observatory image

was obtained on Sept. 7.41 UT.

P. Groot, Harvard-Smithsonian Center for Astrophysics; S. Tingay, Australia Telescope National Facility; A. Udalski, Warsaw University Observatory; and J. Miller, Massachusetts Institute of Technology, report on behalf of a larger collaboration: “Radio observations of the error box of XTE J1650–500 (IAUC 7707) made with the Australia Telescope Compact Array at 1384 MHz on Sept. 7.27–7.58 UT show a new, ∼ 3 mJy source at α = 16^{h}50^{m}01^{s}.00, δ = −49^{o}57′45′′.0 (equinox 2000.0; 3′′ uncertainty), which was 1′.5 away from the RXTE raster-scan position (Markwardt et al., op.cit.). Two 15-min I-band observations of the x-ray error box were obtained with the OGLE 1.3-m telescope at Las Campanas in non-photometric weather on Sept. 7.034 and 7.058. Inspection of the radio position shows the presence of a new object. The new source is 1.8 and 0.8 magnitude fainter than USNO-A2.0 0372-29042342 and 0375-29037362, respectively. Assuming I = 15.3 and 16.5 for these two stars, respectively, the new source is currently at I = 17.2 ± 0.2. The source is barely visible on Second Epoch Red Survey plates taken with the U.K. Schmidt telescope. This source has also been reported by Castro-Tirado et al. (op.cit.) as the possible counterpart. Optical images of the candidate can be found at URLs

http://cfa-www.harvard.edu/~pgroot/XTEJ1650-50DDS.ps and http:
//cfa-www.harvard.edu/~pgroot/XTEJ1650-50OGLE.ps.”

2001 September 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green